OpenStudy (stickman1998):

What is the value of log(81)3

OpenStudy (welshfella):

is 81 the base of the log?

OpenStudy (stickman1998):

Yes

OpenStudy (welshfella):

the log to the base 81 of 3 is the exponent y in the following 81^y = 3

OpenStudy (welshfella):

any ideas what the value of y is?

OpenStudy (stickman1998):

Nope

OpenStudy (catbuck1):

just do log(3)/log(81)

OpenStudy (stickman1998):

It's not 3^1/81 is it?

OpenStudy (welshfella):

Hint: 81 = 3*3*3

OpenStudy (stickman1998):

Is @Catbuck1 correct?

OpenStudy (catbuck1):

i am doing the same thing and i got a 100%

OpenStudy (stickman1998):

I need to put the calculator in rad mode right?

OpenStudy (catbuck1):

naw, just do log(3) divided by log(81) and you got the answer

OpenStudy (welshfella):

No rads are angle values not logs Catbuck1 is wrong

OpenStudy (catbuck1):

how am i wrong? i got a 100% on my quiz with the same question!

OpenStudy (welshfella):

Hint 3 is the cube root of 81 what is the cube root written as a fraction?

OpenStudy (stickman1998):

1/3?

OpenStudy (welshfella):

He's right Sorry 3 is the 4th roots of 81 not the cube.

OpenStudy (catbuck1):

yeah!!!!! i beat a professor and im in eighth grade!

OpenStudy (stickman1998):

so then 1/4^81=3?

OpenStudy (catbuck1):

naw, just 0.25

OpenStudy (welshfella):

log3 / log 81 will give the correct answer 81^ (1/4) = 3 so the log(81)3 = 1/4

OpenStudy (stickman1998):

Alright Ty both :)

OpenStudy (welshfella):

its best to do these problems the way I said because it gives you a better understanding of what logs are. logs are a power ( or exponent).

OpenStudy (welshfella):

yw

OpenStudy (catbuck1):

hey can you help me out? A diner has collected data about customer coffee-drinking habits. They have calculated that P(cream) = 0.5, P(sugar) = 0.6, and P(cream or sugar) = 0.7. Determine the P(cream and sugar).

OpenStudy (stickman1998):

May I ask why you write the exponent on the other side?

OpenStudy (catbuck1):

@welshfella

OpenStudy (welshfella):

lo I dont know what made me think 3 was the cube root of 81!!!!!!

OpenStudy (stickman1998):

lolol

OpenStudy (welshfella):

you dont write 2^x for x squared that means a totally different thing

OpenStudy (welshfella):

professors are human!!

OpenStudy (catbuck1):

A diner has collected data about customer coffee-drinking habits. They have calculated that P(cream) = 0.5, P(sugar) = 0.6, and P(cream or sugar) = 0.7. Determine the P(cream and sugar).? @welshfella